# 03F40 (primary), 03D20

### v2

@@ 4 Termination
Goodstein’s theorem is then proven using the well-foundedness of $\varepsilon(\omega)$, more commonly denoted $\varepsilon_{0}$. For this, we observe that the operation $\left(\begin{smallmatrix}c\\ b\end{smallmatrix}\right)n$ is well-defined even when $c$ is infinite, as it relies on operations available for arbitrary ordinals. In particular, we may take $c=\omega$. The analogue of Lemma 2.1 is then seen to hold, and, specifically, the $\omega$-base-change has the following crucial properties:
@@
Then, $o_{B}(7)=\vartheta(O_{B}(7)\oplus o^{*}_{B}(7))$, where $o^{*}_{B}(7)$ is a countable ordinal large enough to ensure monotonicity of the ordinal assignment by making the maximal coefficient of $O_{B}(n)$ be large enough. [^2] For readers familiar with Buchholz’s fundamental sequences for $\vartheta$ [4], $o^{*}_{B}$ plays a role similar to that of $\vartheta^{*}$. In this particular case, we will have $o^{*}_{B}(7)=o_{B}(6)$, and, in general, $o^{*}_{B}(n)$ will always be either zero or of the form $o_{B}(b^{*})$ for some $b^{*}\in B$. For this to work smoothly, $O_{B}(b)$ will have a slightly different definition when $b\in B$ (see Definition 4.8 below).
@@
Let $b\geq 2$ and $f\colon\mathbb{N}\rightharpoonup{\Omega}$ a partial function with $0,b)\subseteq{\rm dom}(f)$. Then, for $n\in\mathbb{N}$, $O_{f}^{b}(n)$ is a multiple of $\Omega$} if and only if $n$ is a multiple of $b$.
@@
If $r=0$ and $a=1$, note that $e>1$, since $n>b$. Then, $b^{e}-1=b^{e-1}(b-1)+b^{e-1}-1$ and thus $O(b^{e}-1)={\Omega}^{O(e-1)}f(b-1)+O(b^{e-1}-1)$. The induction hypothesis yields O(e1)<O(e)O(e-1)<O(e) and $O(b^{e-1}-1)<O(b^{e-1})$, so since $f(b-1)<{\Omega}$, $O(b^{e}-1)<O(b^{e})$.
@@
where $O_{B}(n)$, $o^{*}_{B}(n)$ are defined according to the following cases.
@@ 1. Case 1
    ($n\geq\min B$). Let $b={\rm base}_{B}(n)$. For $f=o_{B}\upharpoonright[0,b)$, we define an auxiliary operation given by $O^{b}_{B}(x):=O_{f}^{b}(x)$ for all $x\in\mathbb{N}$. Then, consider the following sub-cases.
    1. Case 1.1
@@
Although not strictly necessary, is also convenient to set $O_{B}(n)=n$ and $o^{*}_{B}(n)=0$ if $n<b$, simply to avoid cluttering lemmas with the assumption that $n\geq\max B$.
@@
and $o^{*}_{B}(6)=0$ since $O_{B}(2)+{\Omega}<O_{B}(6)+{\Omega}$. It follows that $o_{B}(6)=\vartheta({\Omega}^{\Omega}+\vartheta({\Omega}^{\Omega}))$.
Next, we see that $O_{B}(7)=O_{B}(6+1)={\Omega}+1$, and in this case $o^{*}_{B}(7)=o_{B}(6)$, so
@@
Meanwhile, both of these ordinals are different from the value we would obtain if we had $4\notin C$: namely, $O^{2}_{C}(4)={\Omega}^{\Omega}$.
@@
Proof 4.12 (sketch).
By induction on $n$. Write $n=_{b}b^{e}x+y$ and observe that either $y=r$ and we can apply Definition 4.4 directly, or else $b<y<S_{B}(b)$ and we can apply the induction hypothesis to $y$.
@@
    establishing the second claim when ${\rm base}_{B}(m)=b$. We moreover note that $O_{B}(m)^{*}$ is of the form $o_{B}(m^{\prime})$ with $m^{\prime}<m$. By the induction hypothesis, $o_{B}(m^{\prime})<o_{B}(n)$, so $o_{B}(m)<o_{B}(n)$ by Proposition 4.3.
@@ 2. Case 2
        Since $O_{B}(m)=O^{b}_{B}(m)$, we have by Lemma 4.11 that
@@
We note that $O_{B}(n)$ is computed differently depending on whether or not $n\in B$. One intuition here is that, since ${\uparrow}_{B}^{C}n=S_{C}({\uparrow}(n-1))$, which is not ‘much larger’ than ${\uparrow}(n-1)$ when $C$ is a good successor of $B$, we should also have $O_{B}(n)$ be not too much larger than $O_{B}(n-1)$. Accordingly, the recipe for the case $n\notin B$ will yield larger values than that for $n\in B$, as made precise by the following.
@@
Proof 4.16.
Since $n$ is a multiple of $b$ by the definition of base hierarchy, $O^{b}_{B}(n)$ and $O^{b}_{B}(n-b)$ are multiples of $\Omega$} by Lemma 4.5. By definition,
@@
$O^{b}_{B}(n-b)+{\Omega}={\Omega}({\alpha}+1)\leq{\Omega}{\beta}=O^{b}_{B}(n).$
@@
The final ingredient in the proof of termination is the preservation of the ordinal interpretation under the base upgrade. Note that if we wish to prove that $o_{B}(m)=o_{C}({{\uparrow}}m)$, we will first need to show that $O_{B}(m)=O_{C}({{\uparrow}}m)$. The following technical lemma will be a first approximation to this.
@@
$O^{b}_{B}(n)={\Omega}^{O^{b}_{B}(e)}o_{B}(a)+O^{b}_{B}(r)={\Omega}^{O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e)}o_{C}({\uparrow}a)+O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r)=O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n).$
We remark that the assumption that $C$ is a good successor for $B$ is not used explicitly, but it ensures that $\uparrow$} is well defined. Lemma 4.18 can then be used in an inductive argument to show that $O_{B}$ is preserved by the upgrade.
@@ 1. Case 1
    ($n\notin B$). We have by Lemma 3.42 that for $d={\rm base}_{C}({\uparrow}n)$, ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$. By Lemma 3.41, ${\uparrow}n\notin C$, so Lemma 4.18 yields
@@ 2. Case 2
    where the second and third equalities are by items 1 and 2 of Lemma 3.10, so in particular, $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba={\uparrow}ba$. By Lemma 3.12, ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba+d$ and ${\rm base}_{C}({\uparrow}(n-1))={\rm Base}_{C}({\uparrow}(n-1))=d$ (where the first equality uses $n-1\notin B$, since $n\in B$). Lemma 4.18 yields $O^{d}_{C}({\uparrow}ba)=O^{b}_{B}(ba)$, so
@@
Proof 4.23.
Let all transformations be from $B$ to $C$. For $n<b$, $o_{B}(n)=o_{C}({\uparrow}n)=n$, so we may assume that $n\geq b$. Proceed by induction on $n$ to show that $O_{B}(n)=O_{C}({\uparrow}n)$ and $o^{*}_{B}(n)=o^{*}_{C}({\uparrow}n)$, from which $o_{B}(n)=o_{C}({\uparrow}n)$ follows by the way we defined $o$. Lemma 4.20 and the induction hypothesis yield $O_{B}(n)=O_{C}({\uparrow}n)$, so it remains to show that $o^{*}_{B}(n)=o^{*}_{C}({\uparrow}n)$.
First, we check that $o^{*}_{B}(n)\leq o^{*}_{C}({{\uparrow}}n)$. The case where $o^{*}_{B}(n)=0$ is immediate, so assume $o^{*}_{B}(n)=o_{B}(b^{*})$ for some $b^{*}\leq b$. By Lemma 4.20,
@@
Since ${\uparrow}b^{*}\in C$ by Lemma 3.41 and ${\uparrow}b^{*}<{\uparrow}n$ by Lemma 3.11, Definition 4.4 yields $o^{*}_{C}({\uparrow}n)=o_{C}(c^{*})$ for some $c^{*}\geq{\uparrow}b^{*}$. Proposition 4.13 and the induction hypothesis then yield $o_{B}(b^{*})=o_{C}({\uparrow}b^{*})\leq o_{C}(c^{*})$. We conclude that $o^{*}_{B}(n)\leq o^{*}_{C}({\uparrow}n)$.
Next, we check that $o^{*}_{C}({{\uparrow}}n)\leq o^{*}_{B}(n)$. As above, we may assume that $o^{*}_{C}({{\uparrow}}n)=o_{C}(c^{*})$ for some $c^{*}\leq{\rm base}_{C}({{\uparrow}}n)$. Let $b^{*}$ be greatest so that $b^{*}\in B$ and ${\uparrow}b^{*}\leq c^{*}$, which exists since ${\uparrow}\min B=\min C\leq c^{*}$. We claim that ${\uparrow}b^{*}=c^{*}$.
If not, since ${\uparrow}b^{*}\leq c^{*}$ by choice of $b^{*}$, we must have that ${\uparrow}b^{*}<c^{*}<{{\uparrow}}n$. By monotonicity (Lemma 3.11), this implies that $b^{*}<n$.
Claim 1.
@@ 2. Case B
    ($m\in B$). Then, $m^{\prime}=m-1\notin B$, since consecutive numbers cannot both be bases as every base divides its successor. We thus have that $b^{*}\leq m^{\prime}<m$, but $b^{*}\in B$, so $b^{*}\neq m^{\prime}$ and thus $b^{*}<m^{\prime}<m\leq S_{B}(b^{*})$. We conclude that ${\rm base}_{B}(m^{\prime})=b^{*}={\rm base}_{B}(m)$. Moreover, since $m\in B$ and $b^{*}$ was chosen to be the maximum element of $B$ with ${\uparrow}b^{*}\leq c^{*}$, we must have that $c^{*}<{\uparrow}m$. By Lemma 3.61, ${\uparrow}m=S_{C}({\uparrow}(m-1))$, but $c^{*}\in C$, so $c^{*}<S_{C}({\uparrow}(m-1))$ can only happen if ${\uparrow}(m-1)\geq c^{*}$. Thus ${\uparrow}m^{\prime}\geq c^{*}$, as claimed.
Let $c={\rm base}_{C}(c^{*})$, and note that $c<c^{*}$, since $\min C\leq{\uparrow}b^{*}<c^{*}$. We thus have that $c^{*}=S_{C}(c)$. Moreover, ${\uparrow}b^{*}\in C$ by Lemma 3.41, so ${\uparrow}b^{*}<c^{*}$ yields ${\uparrow}b^{*}\leq c$.
We claim that $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\geq c^{*}$. Otherwise, $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<c^{*}=S_{C}(c)$ and ${\uparrow}b^{*}\leq c$, and since $m^{\prime}\notin B$, Lemma 3.62 implies that ${\uparrow}m^{\prime}=\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}$ for the least $d\leq c$ with ${\uparrow}b^{*}\leq d$ and $\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<S_{C}(d)$. Since $c$ has these properties, $d\leq c$. Lemma 3.12 would then yield $\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\leq\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<c_{*}$ and thus ${\uparrow}m^{\prime}<c^{*}$, contrary to Claim 1. We conclude that $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\geq c^{*}$, as desired.
@@
$\geq O_{B}(m^{\prime})$ by Proposition 4.13 and Claim 1,
@@
By definition of $m$, either $m=S_{B}(b^{*})<n$ or $m=n$. If $m=S_{B}(b^{*})<n$, we have that ${\uparrow}m<{\uparrow}n$ by Proposition 4.13 and ${\uparrow}m\in C$ by Lemma 3.41. Moreover,
@@
where the second inequality is by choice of $c^{*}$. But $O_{C}({\uparrow}m)\geq O_{C}({\uparrow}n)$ implies that $O_{C}({\uparrow}m)+{\Omega}\geq O_{C}({\uparrow}n)+{\Omega}$, and thus $c^{*}\geq{\uparrow}m$ by definition of $o^{*}_{C}$. Since $m=S_{B}(b^{*})>b^{*}$, this contradicts our choice of $b^{*}$ as the maximum element of $B$ with ${\uparrow}b^{*}\leq c$.
@@
Since $b^{*}\in B$ and $b^{*}<n$, Definition 4.4 yields $o^{*}_{B}(n)=o_{B}(b^{+})$ for some $b^{+}\geq b^{*}$, and Proposition 4.13 and the induction hypothesis then yield $o_{B}(b^{+})\geq o_{B}(b^{*})=o_{C}(c^{*})$. We conclude that $o^{*}_{B}(n)\geq o^{*}_{C}({\uparrow}n)$, and thus equality holds.
@@
where the first inequality is by Proposition 4.13 and the second by Proposition 4.22. It follows that $(o_{i}(\mathbb{G}^{\mathcal{B}}_{i}(m)))_{i<I}$ is a decreasing sequence of ordinals, hence it is finite and we must have $o_{I-1}(\mathbb{G}^{\mathcal{B}}_{I-1}(m))=0$.
5 The Ouroboros Dynamical Hierarchy

### v4

@@ 4. Termination
Goodstein’s theorem is then proven using the well-foundedness of $\varepsilon(\omega)$, more commonly denoted $\varepsilon_{0}$. For this, we observe that $\left(\begin{smallmatrix}c\\ b\end{smallmatrix}\right)n$ is well-defined even when $c$ is infinite, as it relies on operations available for arbitrary ordinals. In particular, we may take $c=\omega$. The analogue of Lemma 2.1 is then seen to hold, and, specifically, the $\omega$-base-change has the following crucial properties:
@@
Then, $o_{B}(7)=\vartheta(O_{B}(7)\oplus o^{*}_{B}(7))$, where $o^{*}_{B}(7)$ is a countable ordinal large enough to ensure monotonicity of the ordinal assignment, by making the maximal coefficient of $O_{B}(n)$ be large enough. [^3] For readers familiar with Buchholz’s fundamental sequences for $\vartheta$ [4], $o^{*}_{B}$ plays a role similar to that of $\vartheta^{*}$. In this particular case, we will have $o^{*}_{B}(7)=o_{B}(6)$, and, in general, $o^{*}_{B}(n)$ will always be either zero or of the form $o_{B}(b^{*})$ for some $b^{*}\in B$. For this to work smoothly, $O_{B}(b)$ will have a slightly different definition when $b\in B$ (see Definition 4.7 below).
@@
Let $b\geq 2$ and $f\colon\mathbb{N}\rightharpoonup{\Omega}$ a partial function with $0,b)\subseteq{\rm dom}(f)$ and such that $f(n)=0$ if and only if $n=0$. Then, for $n\in\mathbb{N}$, $O_{f}^{b}(n)$ is a multiple of $\Omega$} if and only if $n$ is a multiple of $b$.
@@
If $r=0$ and $a=1$, note that $e>1$, since $n>b$. Then, $b^{e}-1=_{b}b^{e-1}(b-1)+b^{e-1}-1$ and thus $O(b^{e}-1)={\Omega}^{O(e-1)}f(b-1)+O(b^{e-1}-1)$. The induction hypothesis yields O(e1)<O(e)O(e-1)<O(e) and $O(b^{e-1}-1)<O(b^{e-1})$, so since $f(b-1)<{\Omega}$, $O(b^{e}-1)<O(b^{e})$. ∎
@@
where $O_{B}(n)$, $o^{*}_{B}(n)$ are defined as follows.
Let $b={\rm base}_{B}(n)$. For $f=o_{B}\upharpoonright[0,b)$, we define an auxiliary operation given by $O^{b}_{B}(x):=O_{f}^{b}(x)$ for all $x\in\mathbb{N}$. Then, consider the following cases.
@@ 2. Case 2
Although not strictly necessary, is also convenient to set $O_{B}(n)=O^{b}_{B}(n)=n$ and $o^{*}_{B}(n)=0$ if $n<b$, simply to avoid cluttering lemmas with the assumption that $n\geq\max B$.
@@
and $o^{*}_{B}(6)=0$, since $O_{B}(2)+{\Omega}<O_{B}(6)+{\Omega}$. It follows that $o_{B}(6)=\vartheta({\Omega}^{\Omega}+\vartheta({\Omega}^{\Omega}))$.
Next, we see that $O_{B}(7)=O_{B}(6+1)={\Omega}+1$, and in this case, $o^{*}_{B}(7)=o_{B}(6)$, so
@@
Meanwhile, both of these ordinals are different from the value we would obtain if we removed $4$ from $C$: namely, $O^{2}_{C}(4)={\Omega}^{\Omega}$.
@@
Proof sketch.
By induction on $n$. Write $n=_{b}b^{e}x+y$ and observe that either $y=r$ and we can apply Definition 4.4 directly, or else $b<y<S_{B}(b)$, and we can apply the induction hypothesis to $y$. ∎
@@ 1. Case 1
    establishing the second claim when ${\rm base}_{B}(m)=b$. We moreover note that $O_{B}(m)^{*}$ is of the form $o_{B}(m^{\prime})$ with $m^{\prime}<m$. By the induction hypothesis, $o_{B}(m^{\prime})<o_{B}(m)$. Similarly, $o^{*}_{B}(m)=o_{B}(b^{*})$ for some $b^{*}<m$ and hence $o^{*}_{B}(m)<o_{B}(n)$. Clearly, $(O_{B}(m)\oplus o_{B}(m))^{*}\leq o_{B}(m^{\prime})\oplus o^{*}_{B}(m)<o_{B}(n)$, where the last inequality uses the fact that $\vartheta(\xi)$ is always additively indecomposable. We conclude that $o_{B}(m)<o_{B}(n)$ by Proposition 4.3.
@@ 2. Case 2
        Since $O_{B}(m)=O^{b}_{B}(m)$, we have by Lemma 4.10 and the induction hypothesis for the first claim that
@@
One intuition for why $O_{B}(n)$ is computed differently depending on whether or not $n\in B$ is that, since ${\uparrow}_{B}^{C}n=S_{C}({\uparrow}(n-1))$, which is not ‘much larger’ than ${\uparrow}(n-1)$ when $C$ is a good successor of $B$, we should also have $O_{B}(n)$ be not too much larger than $O_{B}(n-1)$. Accordingly, the recipe for the case $n\notin B$ will yield larger values than that for $n\in B$, as made precise by the following.
@@
Proof.
Since $n$ is a multiple of $b$ by the definition of a base hierarchy, $O^{b}_{B}(n)$ and $O^{b}_{B}(n-b)$ are multiples of $\Omega$} by Lemma 4.5. By definition,
@@
$O_{B}(n)<O^{b}_{B}(n-b)+{\Omega}={\Omega}({\alpha}+1)\leq{\Omega}{\beta}=O^{b}_{B}(n).\qed$
@@
The final ingredient in the proof of termination is the preservation of the ordinal interpretation under the upgrade. Note that if we wish to prove that $o_{B}(m)=o_{C}({{\uparrow}}m)$, we will first need to show that $O_{B}(m)=O_{C}({{\uparrow}}m)$. The following technical lemma will be a first approximation to this.
@@
$O^{b}_{B}(n)={\Omega}^{O^{b}_{B}(e)}o_{B}(a)+O^{b}_{B}(r)={\Omega}^{O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle e)}o_{C}({\uparrow}a)+O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle r)=O^{c}_{C}(\langle\begin{smallmatrix}c\\ b\end{smallmatrix}\rangle n).\qed$
We remark that the assumption that $C$ is a good successor for $B$ is not used explicitly in Lemma 4.14, but it ensures that $\uparrow$} is well defined. This lemma can then be used in an inductive argument to show that $O_{B}$ is preserved by the upgrade.
@@ 1. Case 1
    ($n\notin B$). We have by Lemma 3.3(2) that for $d={\rm base}_{C}({\uparrow}n)$, ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$. By Lemma 3.3(1), ${\uparrow}n\notin C$, so Lemma 4.14 yields
@@ 2. Case 2
    where the second and third equalities are by items (1) and (2) of Lemma 3.6, so in particular, $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba={\uparrow}ba$. By Lemma 3.7, ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle ba+d$ and ${\rm base}_{C}({\uparrow}(n-1))={\rm Base}_{C}({\uparrow}(n-1))=d$ (where the first equality uses $n-1\notin B$, since $n\in B$). Lemma 4.14 yields $O^{d}_{C}({\uparrow}ba)=O^{b}_{B}(ba)$, so
@@
Proof.
Let all transformations be from $B$ to $C$. For $n<\min B$, $o_{B}(n)=o_{C}({\uparrow}n)=n$, so we may assume that $n\geq\min B$. Proceeding by induction on $n$, Lemma 4.15 and the induction hypothesis yield $O_{B}(n)=O_{C}({\uparrow}n)$. It remains to show that $o^{*}_{B}(n)=o^{*}_{C}({\uparrow}n)$, from which we can conclude that
$o_{B}(n)=\vartheta(O_{B}(n)\oplus o^{*}_{B}(n))=\vartheta(O_{C}({\uparrow}n)\oplus o^{*}_{C}({\uparrow}n))=o_{C}({\uparrow}n),$
as desired.
First, we check that $o^{*}_{B}(n)\leq o^{*}_{C}({{\uparrow}}n)$. The case where $o^{*}_{B}(n)=0$ is immediate, so assume $o^{*}_{B}(n)=o_{B}(b^{*})$ for some $b^{*}\in B$. By Lemma 4.15,
@@
Since ${\uparrow}b^{*}\in C$ by Lemma 3.3(1) and ${\uparrow}b^{*}<{\uparrow}n$ by Lemma 3.1(1), Definition 4.4 yields $o^{*}_{C}({\uparrow}n)=o_{C}(c^{*})$ for some $c^{*}\geq{\uparrow}b^{*}$. Proposition 4.11 and the induction hypothesis then yield $o_{B}(b^{*})=o_{C}({\uparrow}b^{*})\leq o_{C}(c^{*})$. We conclude that $o^{*}_{B}(n)\leq o^{*}_{C}({\uparrow}n)$.
Next, we check that $o^{*}_{C}({{\uparrow}}n)\leq o^{*}_{B}(n)$. As above, we may assume that $o^{*}_{C}({{\uparrow}}n)=o_{C}(c^{*})$ for some $c^{*}\in C$. Let $b^{*}$ be greatest so that $b^{*}\in B$ and ${\uparrow}b^{*}\leq c^{*}$, which exists since ${\uparrow}\min B=\min C\leq c^{*}$. We claim that ${\uparrow}b^{*}=c^{*}$.
If not, since ${\uparrow}b^{*}\leq c^{*}$ by choice of $b^{*}$, we must have that ${\uparrow}b^{*}<c^{*}<{{\uparrow}}n$. By monotonicity (Lemma 3.1(1)), this implies that $b^{*}<n$.
Claim 4.17.
@@ 2. Case B
    ($m\in B$). Then, $m=S_{B}(b^{*})$ and $m^{\prime}=m-1\notin B$, since consecutive numbers cannot both be bases, given that every base divides its successor. We thus have that $b^{*}\leq m^{\prime}<m$, but $b^{*}\in B$, so $b^{*}\neq m^{\prime}$ and thus $b^{*}<m^{\prime}<m=S_{B}(b^{*})$. We conclude that ${\rm base}_{B}(m^{\prime})=b^{*}={\rm base}_{B}(m)$. Moreover, since $m\in B$ and $b^{*}$ was chosen to be the maximum element of $B$ with ${\uparrow}b^{*}\leq c^{*}$, we must have that $c^{*}<{\uparrow}m$. By Lemma 3.4(1), ${\uparrow}m=S_{C}({\uparrow}(m-1))$, but $c^{*}\in C$, so $c^{*}<S_{C}({\uparrow}(m-1))$ can only happen if ${\uparrow}(m-1)\geq c^{*}$. Thus ${\uparrow}m^{\prime}\geq c^{*}$, as claimed.∎
Let $c={\rm base}_{C}(c^{*})$, and note that $c<c^{*}$, since $\min C\leq{\uparrow}b^{*}<c^{*}$. We thus have that $c^{*}=S_{C}(c)$. Moreover, ${\uparrow}b^{*}\in C$ by Lemma 3.3(1), so ${\uparrow}b^{*}<c^{*}$ yields ${\uparrow}b^{*}\leq c$.
We claim that $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\geq c^{*}$. Otherwise, $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<c^{*}=S_{C}(c)$ and ${\uparrow}b^{*}\leq c$, and since $m^{\prime}\notin B$, Lemma 3.4(2) implies that ${\uparrow}m^{\prime}=\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}$ for the least $d\leq c$ with ${\uparrow}b^{*}\leq d$ and $\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<S_{C}(d)$. Since $c$ has these properties, $d\leq c$. Lemma 3.1(3) would then yield $\langle\begin{smallmatrix}d\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\leq\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}<c_{*}$ and thus ${\uparrow}m^{\prime}<c^{*}$, contrary to Claim 4.17. We conclude that $\langle\begin{smallmatrix}c\\ b^{*}\end{smallmatrix}\rangle m^{\prime}\geq c^{*}$, as desired.
@@
$\geq O_{B}(m^{\prime})$ by Proposition 4.11 and Claim 4.17,
@@
By definition of $m$, either $m=S_{B}(b^{*})<n$ or $m=n$. If $m=S_{B}(b^{*})<n$, we have that ${\uparrow}m<{\uparrow}n$ by Proposition 4.11 and ${\uparrow}m\in C$ by Lemma 3.3(1). Moreover,
@@
where the second inequality is by choice of $c^{*}$. But $O_{C}({\uparrow}m)\geq O_{C}({\uparrow}n)$ implies that $O_{C}({\uparrow}m)+{\Omega}\geq O_{C}({\uparrow}n)+{\Omega}$, and thus $c^{*}\geq{\uparrow}m$ by definition of $o^{*}_{C}$. But Lemma 3.1(1) and Claim 4.17 yield ${\uparrow}m>{\uparrow}(m-1)={\uparrow}m^{\prime}\geq c^{*}$, a contradiction.
@@
Since $b^{*}\in B$ and $b^{*}<n$, Definition 4.4 yields $o^{*}_{B}(n)=o_{B}(b^{+})$ for some $b^{+}\geq b^{*}$, and Proposition 4.11 and the induction hypothesis then yield $o_{B}(b^{+})\geq o_{B}(b^{*})=o_{C}(c^{*})$. We conclude that $o^{*}_{B}(n)\geq o^{*}_{C}({\uparrow}n)$, and thus equality holds. ∎
@@
where the first inequality is by Proposition 4.11 and the second by Proposition 4.16. It follows that $(o_{i}(\mathbb{G}^{\mathcal{B}}_{i}(m)))_{i<I}$ is a decreasing sequence of ordinals, hence it is finite, and we must have $o_{I-1}(\mathbb{G}^{\mathcal{B}}_{I-1}(m))=0$. ∎
5. The Greedy Dynamical Hierarchy